# AdaSCALE: Adaptive Scaling for OOD Detection

### v1

@@ Appendix A Notations
Appendix B Additional studies
Image Perturbation.
A sufficiently small perturbation, as discussed in Sec. 4.3, is used to make distinctive observations in ID vs OOD. We now systematically investigate the impact of the extent and nature of perturbation on pre-determining OOD likelihood which is in-turn responsible for OOD detection. We present the results in Table 9. It is clearly evident that perturbing regions with trivial intensity value leads to better OOD detection. We can infer that with such perturbation, acts as a more reliable proxy for OOD likelihood. Another key takeaway is that randomly perturbing small regions can also lead to strong performance. This finding highlights a more efficient approach to OOD detection, relying solely on minor random perturbations rather than requiring a predetermination of trivial pixels. Furthermore, the sensitivity study of presented at Table 10 suggests the optimal value of to be around 0.5.
Region type OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD Random 61.73 / 78.15 19.44 / 95.74 83.19 / 48.59 53.45 / 74.10 59.97 / 78.67 18.14 / 96.06 81.92 / 49.19 52.35 / 74.84 60.27 / 78.02 18.45 / 96.00 82.07 / 48.62 52.84 / 74.70 62.81 / 76.27 19.95 / 95.70 83.40 / 46.55 54.94 / 73.42 Trivial 61.77 / 78.29 19.28 / 95.77 82.92 / 48.86 53.34 / 74.24 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89 60.24 / 78.17 17.94 / 96.08 82.19 / 48.59 52.59 / 74.77 66.43 / 74.10 21.58 / 95.29 85.26 / 44.86 56.56 / 72.82 Significant 69.43 / 75.13 22.62 / 95.17 85.59 / 48.32 53.76 / 75.23 67.31 / 75.78 21.24 / 95.44 85.07 / 48.40 53.26 / 75.63 65.65 / 76.20 20.39 / 95.62 84.36 / 48.33 53.17 / 75.55 64.54 / 75.39 20.48 / 95.61 83.78 / 47.07 54.11 / 74.54 All 67.37 / 73.08 22.93 / 95.01 85.66 / 44.00 57.80 / 72.19
Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC 0.1 63.76 77.50 19.26 95.85 0.5 58.97 78.98 17.84 96.14 1.0 61.60 76.96 19.31 95.84
Latency.
AdaSCALE incurs extra computational overhead in comparison to its preceding works [7, 54, 8]. It uses an extra forward pass to obtain perturbed activation . On top of that, top-k operations (time complexity: , where D is dimension of activation space) on and are employed for pre-determining OOD likelihood. Furthermore, comparing variable percentile to fixed percentile for scaling in Table 11 across 10000 trials, we observe variable percentile inducing relatively higher latency. The latency ratio decreases as activation space with a higher dimension is considered. All experiments are performed using NVIDIA RTX 4090 GPUs.
Fixed percentile (SCALE) 33 40 45 48 54 Variable percentile (AdaSCALE) 152 149 155 152 164 Latency ratio (AdaSCALE / SCALE) 4.66 3.76 3.42 3.14 3.02
Appendix C Hyperparameters
All hyperparameters are determined with respect to the AUROC metric using automatic parameter search of OpenOOD [57, 61]. Although AdaSCALE may appear to require numerous hyperparameters, our findings indicate that setting to consistently yields near-optimal performance across all setups. Consequently, it can be inferred that only the hyperparameters and need to be appropriately tuned for any new architecture for near-optimal performance.
| Dataset | Network | Hyperparameters | ||||||
| CIFAR-10 | WideResNet-28-10 | 60 | 95 | 10 | 0.5 | |||
| DenseNet-101 | 65 | 90 | 10 | 0.5 | ||||
| CIFAR-100 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
| DenseNet-101 | 70 | 80 | 10 | 0.5 | ||||
| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
| ResNet-101 | 80 | 85 | 10 | 0.5 | ||||
| RegNet-Y-16 | 60 | 90 | 10 | 0.5 | ||||
| ResNeXt-50 | 80 | 85 | 10 | 0.5 | ||||
| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
| Vit-B-16 | 60 | 85 | 10 | 0.5 | ||||
| Swin-B | 90 | 99 | 10 | 0.5 | ||||
| Dataset | Network | Hyperparameters | ||||||
| CIFAR-10 | WideResNet-28-10 | 60 | 85 | 10 | 0.5 | |||
| DenseNet-101 | 70 | 85 | 10 | 0.5 | ||||
| CIFAR-100 | WideResNet-28-10 | 60 | 80 | 10 | 0.5 | |||
| DenseNet-101 | 65 | 75 | 10 | 0.5 | ||||
| ImageNet-1k | ResNet-50 | 80 | 85 | 10 | 0.5 | |||
| ResNet-101 | 70 | 80 | 10 | 0.5 | ||||
| RegNet-Y-16 | 60 | 85 | 10 | 0.5 | ||||
| ResNeXt-50 | 70 | 80 | 10 | 0.5 | ||||
| DenseNet-201 | 90 | 95 | 10 | 0.5 | ||||
| EfficientNetV2-L | 60 | 99 | 10 | 0.5 | ||||
| Vit-B-16 | 75 | 85 | 10 | 0.5 | ||||
| Swin-B | 90 | 99 | 10 | 0.5 | ||||
Appendix D CIFAR-results
D.1 WRN-28-10
@@ D.2 DenseNet-101
Appendix E ImageNet-1k results
E.1 near-OOD detection
@@ Appendix E ImageNet-1k results
E.3 Full-Spectrum near-OOD detection

### v2

@@ Appendix A Notations
Appendix B Algorithm
The algorithm for computing adaptive scaling factor is provided in Algorithm˜1.
Appendix C Additional observation in activation space.
Figure˜5 shows perturbed activations are, on average, higher for ID samples than for OOD samples.
Appendix D Additional Experiments and Analysis
D.1 Superiority of the Proposed OODness Metric ()
To validate our choice of the OODness metric, we conducted an extensive ablation study comparing the performance of our proposed metric, against alone. The results, summarized in Table 12, demonstrate that using provides a consistent and significant performance improvement over using alone. This advantage holds across all eight diverse architectures.
| Category | Metric | ResNet-50 | ResNet-101 | RegNet-Y-16 | ResNeXt-50 | DenseNet-201 | EfficientNetV2-L | ViT-B-16 | Swin-B | Average |
| near-OOD | 59.65/77.82 | 56.67/81.65 | 60.63/86.00 | 63.99/79.34 | 68.65/78.03 | 73.18/80.17 | 82.61/69.31 | 75.85/72.26 | 68.90/78.07 | |
| 58.98/78.98 | 57.96/81.68 | 47.91/89.18 | 64.14/79.96 | 61.28/79.66 | 53.78/86.94 | 71.87/73.14 | 73.41/74.48 | 61.17/80.50 | ||
| far-OOD | 19.82/95.72 | 20.32/95.50 | 31.53/93.94 | 26.87/94.32 | 35.55/92.19 | 66.20/84.60 | 77.52/76.39 | 63.20/80.78 | 42.63/89.18 | |
| 17.84/96.14 | 18.51/95.95 | 21.37/95.84 | 22.08/95.24 | 28.01/93.23 | 37.61/91.48 | 47.63/86.83 | 47.81/87.14 | 30.11/92.73 |
For instance, if an OOD sample gets a very high score, it will produce a very large scaling factor. This can amplify the logits so much that it washes out the subtle but important information contained within the values of the logits themselves. By adding the opposing term, we "tame" these potentially overconfident scaling factors. This ensures a healthier balance where both the feature-level instability (captured by ) and the original logit distribution contribute to the final OOD score.
Illustrative Example: A simple toy example demonstrates why this balancing is critical.
Logits: Let and .
Case 1: Overconfident Scaling (using only )
Let’s say this yields scaling factors and . Assume a threshold of 25, where energy > 25 is ID and energy < 25 is OOD (We deal with magnitude for simplicity).
Here, , leading to an OOD detection failure.
Case 2: Balanced Scaling (using to temper the scaling)
This leads to less extreme scaling, e.g., and . Assume a threshold of 5, where energy > 5.5 is ID and energy < 5.5 is OOD (We deal with magnitude for simplicity).
Here, , leading to a successful OOD detection.
D.2 Comparison with Distance-Based OOD Detection Methods
To situate AdaSCALE within the broader landscape of OOD detection techniques, we compare its performance against representative distance-based methods, specifically NNGuide and MDS. The experiments were conducted on the large-scale ImageNet-1k benchmark using a ResNet-50 backbone. As shown in Table 13, AdaSCALE demonstrates a significant performance advantage over these methods, particularly in the challenging near-OOD setting.
| Method | Near-OOD | Far-OOD | ||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| NNGuide | 72.46 | 68.09 | 21.40 | 95.33 |
| MDS | 76.27 | 64.53 | 70.82 | 66.22 |
| AdaSCALE-A | 58.97 | 78.98 | 17.84 | 96.14 |
D.3 Analysis on Adversarially Trained Models
We investigate the effectiveness of our approach on adversarially trained models to test its robustness. This analysis is twofold: first, we verify that the core assumption of our method holds, and second, we evaluate its OOD detection performance.
D.3.1 Activation Sensitivity in Adversarially Trained Networks
The effectiveness of our method relies on the activation shift ratio () being greater than 1 for OOD samples compared to ID samples. We measured this ratio by taking the top-1% activation shifts on a ResNet-50 model adversarially trained on ImageNet-1k. The results in Table 14 confirm that the ratio remains consistently greater than 1 across various OOD datasets. This demonstrates that OOD samples still produce a larger activation shift than ID samples, even in adversarially robust models, validating our method’s applicability.
| OOD Dataset | () | () |
| SSB-hard | 1.78 | 1.73 |
| NINCO | 1.29 | 1.28 |
| ImageNet-O | 1.69 | 1.72 |
| OpenImage-O | 1.48 | 1.45 |
| iNaturalist | 1.14 | 1.14 |
| Textures | 1.33 | 1.18 |
| Places | 1.30 | 1.16 |
D.3.2 Performance of AdaSCALE on Adversarially Trained Networks
We evaluate AdaSCALE’s OOD detection performance on these adversarially trained models using the exact same hyperparameters as for the standard ResNet-50. The results in Table 15 show that AdaSCALE consistently outperforms baseline methods. This highlights the robustness of our approach AdaSCALE, which remains superior even when applied to models with different training schemes without any specific hyperparameter tuning.
| Training | Method | Near-OOD | Far-OOD | ||
| FPR@95 | AUROC | FPR@95 | AUROC | ||
| SCALE | 68.54 | 75.43 | 25.95 | 94.40 | |
| OptFS | 70.36 | 72.97 | 28.60 | 93.24 | |
| AdaSCALE-A | 60.73 | 80.84 | 19.90 | 95.74 | |
| SCALE | 66.99 | 76.10 | 24.77 | 94.40 | |
| OptFS | 70.11 | 73.51 | 28.32 | 93.42 | |
| AdaSCALE-A | 60.21 | 80.68 | 21.13 | 95.35 | |
D.4 Robustness to Real-World Image Corruptions
To evaluate the practical robustness of our method, we assess its performance when input images are subjected to common real-world corruptions. We benchmark AdaSCALE against SCALE and OptFS on the ImageNet-1k dataset using a ResNet-50, applying Gaussian blur and JPEG compression to the input images. The results, presented without any hyperparameter re-tuning, show that AdaSCALE’s performance advantage is maintained even under these challenging conditions.
| Method | Near-OOD | Far-OOD | ||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| SCALE | 72.68 | 70.15 | 29.65 | 92.72 |
| OptFS | 71.84 | 69.66 | 30.27 | 92.24 |
| AdaSCALE-A | 67.53 | 75.18 | 28.11 | 93.31 |
| Method | Near-OOD | Far-OOD | ||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| SCALE | 69.73 | 71.84 | 27.46 | 93.56 |
| OptFS | 69.43 | 71.12 | 28.34 | 92.95 |
| AdaSCALE-A | 61.24 | 77.59 | 25.87 | 94.06 |
Appendix E Sensitivity study
E.1 Sensitivity study of
The sensitivity study of presented at Table˜18 suggests the optimal value of to be around 0.5.
| Near-OOD | Far-OOD | |||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| 0.1 | 63.76 | 77.50 | 19.26 | 95.85 |
| 0.5 | 58.97 | 78.98 | 17.84 | 96.14 |
| 1.0 | 61.60 | 76.96 | 19.31 | 95.84 |
E.2 Sensitivity study of
The sensitivity study of presented at Table˜18 suggests the optimal value of to be around 10.
| Near-OOD | Far-OOD | |||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| 0.1 | 70.07 | 74.12 | 21.40 | 95.33 |
| 1 | 67.49 | 75.45 | 20.41 | 95.54 |
| 10 | 58.97 | 78.98 | 17.84 | 96.14 |
| 100 | 59.03 | 78.10 | 19.39 | 95.80 |
E.3 Sensitivity study of
The sensitivity study of presented at Table˜20 suggests the optimal value of to be around 1%.
| Near-OOD | Far-OOD | |||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| 1% | 58.97 | 78.98 | 17.84 | 96.14 |
| 5% | 60.36 | 77.74 | 19.95 | 95.63 |
| 10% | 60.40 | 77.89 | 19.72 | 95.63 |
| 50% | 60.34 | 77.25 | 20.90 | 95.25 |
| 100% | 60.99 | 76.49 | 21.88 | 94.96 |
E.4 Sensitivity study of
The sensitivity study of presented at Table˜21 suggests the optimal value of to be around 5%.
| Near-OOD | Far-OOD | |||
| FPR@95 | AUROC | FPR@95 | AUROC | |
| 1% | 59.08 | 78.56 | 18.35 | 96.02 |
| 5% | 58.97 | 78.98 | 17.84 | 96.14 |
| 10% | 59.41 | 78.65 | 18.33 | 95.96 |
| 100% | 62.43 | 75.44 | 23.05 | 94.69 |
E.5 Sensitivity study of
We present the complete results of image perturbation study (FPR@95 / AUROC ) in Table˜22.
| Pixel type | OOD Detection | FS-OOD Detection | |||
| Near-OOD | Far-OOD | Near-OOD | Far-OOD | ||
| Random | 61.73 / 78.15 | 19.44 / 95.74 | 83.19 / 48.59 | 53.45 / 74.10 | |
| 59.97 / 78.67 | 18.14 / 96.06 | 81.92 / 49.19 | 52.35 / 74.84 | ||
| 60.27 / 78.02 | 18.45 / 96.00 | 82.07 / 48.62 | 52.84 / 74.70 | ||
| 62.81 / 76.27 | 19.95 / 95.70 | 83.40 / 46.55 | 54.94 / 73.42 | ||
| Trivial | 61.77 / 78.29 | 19.28 / 95.77 | 82.92 / 48.86 | 53.34 / 74.24 | |
| 58.97 / 78.98 | 17.84 / 96.14 | 81.52 / 49.35 | 52.33 / 74.89 | ||
| 60.24 / 78.17 | 17.94 / 96.08 | 82.19 / 48.59 | 52.59 / 74.77 | ||
| 66.43 / 74.10 | 21.58 / 95.29 | 85.26 / 44.86 | 56.56 / 72.82 | ||
| Salient | 69.43 / 75.13 | 22.62 / 95.17 | 85.59 / 48.32 | 53.76 / 75.23 | |
| 67.31 / 75.78 | 21.24 / 95.44 | 85.07 / 48.40 | 53.26 / 75.63 | ||
| 65.65 / 76.20 | 20.39 / 95.62 | 84.36 / 48.33 | 53.17 / 75.55 | ||
| 64.54 / 75.39 | 20.48 / 95.61 | 83.78 / 47.07 | 54.11 / 74.54 | ||
| All | 67.37 / 73.08 | 22.93 / 95.01 | 85.66 / 44.00 | 57.80 / 72.19 | |
E.6 Sensitivity study of
We present sensitivity study (FPR@95 / AUROC ) of , setting to 85 in the Table˜23.
| 70 | 75 | 80 | 85 | |
| Near-OOD | 62.54 / 79.60 | 58.57 / 81.68 | 58.98 / 78.98 | 67.76 / 74.20 |
| Far-OOD | 21.62 / 95.19 | 19.86 / 95.62 | 17.84 / 96.14 | 21.44 / 95.39 |
E.7 Joint sensitivity study
We present joint sensitivity study in terms of near-OOD detection among three hyperparameters (, below:
| \ | 0.1 | 1 | 10 | 100 |
| 1% | 70.07 / 74.12 | 67.49 / 75.45 | 58.97 / 78.98 | 59.03 / 78.10 |
| 5% | 69.53 / 74.33 | 63.87 / 76.84 | 60.36 / 77.74 | 61.12 / 77.22 |
| 10% | 69.34 / 74.50 | 62.63 / 77.33 | 60.40 / 77.89 | 61.55 / 76.92 |
| 50% | 67.49 / 75.32 | 61.26 / 77.26 | 60.34 / 77.25 | 62.05 / 76.52 |
| 100% | 65.95 / 75.77 | 61.18 / 76.97 | 62.17 / 76.39 | 62.43 / 76.31 |
The results in Table˜24 clearly indicate that the hyperparameter exerts a more significant influence on the final OOD detection performance. Furthermore, a key relationship emerges: for optimal performance, a decrease in must be compensated by an increase in (and consequently, a higher ). This is explained by the formulation , where a smaller necessitates a larger value to ensure the term dominates the regularizer .
| \ | 85 | 90 | 95 |
| 0.1 | 70.07 / 74.12 | 87.45 / 66.06 | 91.81 / 56.09 |
| 1 | 67.49 / 75.45 | 84.31 / 69.32 | 89.41 / 60.11 |
| 10 | 58.97 / 78.98 | 69.47 / 79.68 | 78.77 / 76.26 |
| 100 | 59.03 / 78.10 | 63.19 / 79.67 | 71.43 / 78.51 |
In Table˜25, when is set to 0.1 with , does not dominate over term. As a result, it leads to scaling based on the regularization / balancing term instead of "predetermined" actual OOD detection signal "activation shift at peak activations". Such value of (e.g., ) invalidates the proposed hypothesis.
From the results in Table˜26, it can be observed that activation shift with all activations considered () still proves to be a useful OOD detection signal.
Furthermore, previous works (ASH, SCALE) observed that, particularly for ResNet-50 architecture, the performance drops after the percentile value of 85%. Similar observations can also be made in AdaSCALE from the results in Tables˜25 and 26.
| \ | 85 | 90 | 95 |
| 1% | 58.97 / 78.98 | 69.49 / 79.67 | 78.77 / 76.26 |
| 5% | 60.36 / 77.74 | 68.30 / 78.23 | 78.19 / 75.58 |
| 10% | 60.40 / 77.89 | 68.46 / 77.46 | 78.90 / 74.70 |
| 50% | 60.34 / 77.25 | 69.88 / 76.39 | 80.46 / 73.17 |
| 100% | 62.17 / 76.39 | 70.13 / 75.93 | 81.22 / 72.32 |
Appendix F ISH regularization:
Apart from enhancing the prior postprocessor ASH (Djurisic et al., 2023), SCALE (Xu et al., 2015) introduces a training regularization to emphasize samples with more distinct ID characteristics. We assess the performance (FPR@95 / AUROC ) of each method in ResNet-50 and ResNet-101 model following this regularization in Table˜27. The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by and in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively. Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in full-spectrum setting. Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%. These findings also generalize to ResNet-101 network.
| Method | OOD Detection | FS-OOD Detection | ||
| Near-OOD | Far-OOD | Near-OOD | Far-OOD | |
| ResNet-50 | ||||
| MSP | 74.07 / 62.16 | 51.13 / 84.64 | 87.52 / 40.36 | 74.41 / 61.52 |
| MLS | 74.38 / 66.43 | 41.57 / 88.90 | 88.89 / 39.49 | 71.53 / 61.69 |
| EBO | 74.68 / 66.46 | 41.85 / 88.83 | 89.05 / 39.18 | 71.77 / 61.11 |
| ReAct | 71.98 / 70.81 | 28.76 / 93.49 | 87.78 / 43.88 | 61.87 / 71.64 |
| ASH | 67.99 / 73.46 | 23.88 / 94.67 | 85.74 / 45.29 | 57.81 / 72.81 |
| SCALE | 65.68 / 76.41 | 20.77 / 95.62 | 84.31 / 48.40 | 54.48 / 74.79 |
| BFAct | 71.59 / 70.85 | 28.38 / 93.50 | 87.51 / 43.95 | 61.39 / 71.43 |
| LTS | 66.32 / 75.03 | 22.07 / 95.28 | 85.08 / 46.16 | 57.11 / 73.06 |
| OptFS | 67.71 / 73.03 | 24.65 / 94.18 | 85.38 / 45.91 | 57.09 / 72.95 |
| AdaSCALE-A | 57.43 / 80.86 | 16.52 / 96.46 | 80.85 / 51.24 | 51.57 / 75.63 |
| AdaSCALE-L | 56.83 / 80.81 | 17.62 / 96.22 | 80.97 / 50.59 | 53.43 / 74.62 |
| ResNet-101 | ||||
| MSP | 71.39 / 68.31 | 51.00 / 84.81 | 85.70 / 45.72 | 73.69 / 62.79 |
| MLS | 72.32 / 71.94 | 41.04 / 88.99 | 87.39 / 44.88 | 69.94 / 63.23 |
| EBO | 72.78 / 71.92 | 41.45 / 88.87 | 87.64 / 44.66 | 70.23 / 62.65 |
| ReAct | 67.74 / 75.74 | 28.53 / 93.47 | 85.40 / 48.83 | 60.50 / 72.18 |
| ASH | 66.03 / 77.79 | 25.21 / 94.43 | 83.95 / 50.57 | 56.91 / 73.37 |
| SCALE | 64.30 / 78.98 | 23.09 / 94.95 | 83.14 / 51.05 | 55.88 / 73.27 |
| BFAct | 67.53 / 75.86 | 28.32 / 93.40 | 85.11 / 48.96 | 60.12 / 71.83 |
| LTS | 66.32 / 75.03 | 22.07 / 95.28 | 85.08 / 46.16 | 57.11 / 73.06 |
| OptFS | 67.71 / 73.03 | 24.65 / 94.18 | 85.38 / 45.91 | 57.09 / 72.95 |
| AdaSCALE-A | 54.66 / 83.52 | 16.81 / 96.32 | 78.52 / 55.19 | 49.92 / 76.20 |
| AdaSCALE-L | 53.91 / 83.49 | 17.55 / 96.15 | 78.63 / 54.60 | 51.47 / 75.47 |
Appendix G Hyperparameters
| Dataset | Network | ||
| CIFAR-10 | WideResNet-28-10 | 60 | 95 |
| DenseNet-101 | 65 | 90 | |
| CIFAR-100 | WideResNet-28-10 | 60 | 85 |
| DenseNet-101 | 70 | 80 | |
| ImageNet-1k | ResNet-50 | 80 | 85 |
| ResNet-101 | 80 | 85 | |
| RegNet-Y-16 | 60 | 90 | |
| ResNeXt-50 | 80 | 85 | |
| DenseNet-201 | 90 | 95 | |
| EfficientNetV2-L | 60 | 99 | |
| Vit-B-16 | 60 | 85 | |
| Swin-B | 90 | 99 |
| Dataset | Network | ||
| CIFAR-10 | WideResNet-28-10 | 60 | 85 |
| DenseNet-101 | 70 | 85 | |
| CIFAR-100 | WideResNet-28-10 | 60 | 80 |
| DenseNet-101 | 65 | 75 | |
| ImageNet-1k | ResNet-50 | 80 | 85 |
| ResNet-101 | 70 | 80 | |
| RegNet-Y-16 | 60 | 85 | |
| ResNeXt-50 | 70 | 80 | |
| DenseNet-201 | 90 | 95 | |
| EfficientNetV2-L | 60 | 99 | |
| Vit-B-16 | 75 | 85 | |
| Swin-B | 90 | 99 |
Appendix H CIFAR-results
H.1 WRN-28-10
@@ H.2 DenseNet-101
Appendix I ImageNet-1k results
I.1 near-OOD detection
@@ Appendix I ImageNet-1k results
I.3 Full-Spectrum near-OOD detection